The probability of drawing a white marble is .46
What is the probability of a person going into the NFL?
It is pretty close to 0. It is estimated as the number of people in the NFL divided by the total population of the world!
The probability is 0.625
Who would most likely result in an entrepreneur's improving the probability of his or her business?
An entrepreneur improving his or her business in extremely unlikely to result in some person coming into being - unless the entrepreneur is so happy with his/her financial success that he/she decides to make a baby!
What is the probability of spinning red on the first spinner and blue on the second spinner?
The answer will depend on the patter of colours on the two spinners.
What is the definition for probability sample?
A probability sample is one in which each member of the population has the same probability of being included. An alternative and equivalent definition is that it is a sample such that the probability of selecting that particular sample is the same for all samples of that size which could be drawn from the population.
The Answers community requires more information for this question. Please edit your question to include more context. For example, telling us a fair what, without SHOUTING might me a good start.
It is 0.375
It would help if the number of blue disks - or a letter to represent that number - were given.
What does it mean for an event to have a probability of 1?
Probability is on a scale of 1 to 0 and 1 is that .an event will certainly happen whereas 0 is that an event will not happen
It is 1 - prob(it does not rain today and tomorrow)= 1 - prob(it does not rain today)*prob(it does not rain tomorrow)
= 1 - [1- prob(it does rain today)]*[1 - prob(it does rain tomorrow)]
= 1 - [1 - 0.8]*[1 - 0.5]
= 1 - 0.2*0.5
= 1 - 0.1 = 0.9
The probability is 0. One card cannot be a club and a spade!
What is the probability that the spinner will land 5?
The answer depends on the shape of the spinner and the numbers on it.
What is the probability that sack will draw the letter T from the bag?
The answer depends on which letters are in the bag!
What is the probability in percent that a single coin toss will result in heads or tails?
It is 100%. The coin will result in heads or tails since there are no other possible outcomes.
The answer will depend on what the disease is.
How do i find sample standard deviation from population standard deviation?
If the population standard deviation is sigma, then the estimate for the sample standard error for a sample of size n, is s = sigma*sqrt[n/(n-1)]
How many 7 card hands are possible with a 13 cards deck?
If the cards are all different then there are 13C7 = 1716 different hands.